OpenStudy (anonymous):

(Maybe I should change my profile to an attractive girl so I can get this answered =P) You toss a rock up vertically at an initial speed of 30 feet per second and release it at an initial height of 6 feet. The rock will remain in the air for (answer)seconds. It will reach a maximum height of (answer) feet after (answer) seconds. Note: Ignore air resistance.

OpenStudy (phi):

this sounds like physics

OpenStudy (anonymous):

It's for a Calculus class.

OpenStudy (phi):

what formula or equation are you using for this one?

OpenStudy (anonymous):

I don't know, since he hasn't really gone over an equation for this, which is why I'm wondering it's even on my homework. I know that I have to take gravity in to account, which I believe in ft/s is generally 32. Since the rock is already at 6 ft, I know that will be factored in as well. At one point the rock will stop, and gravity will send it back down, so the 2nd velocity is going to be 32ft/s, and the initial velocity is 30 ft/s.

OpenStudy (phi):

g= 32 ft/s^2

OpenStudy (cwrw238):

you can use the equations for constant acceleration which are derived using calculus

OpenStudy (phi):

we could solve for v0 - gt = 0 to find the time to reach 0 velocity

OpenStudy (anonymous):

so, 30 - 32(6)^2?

OpenStudy (anonymous):

wait, ignore that 6.

OpenStudy (anonymous):

t = -30/32?

OpenStudy (anonymous):

30/32*

OpenStudy (phi):

yes, it takes that long to reach the peak. distance will be d= x0 + v0 * t - g t^2 that will be the max height.

OpenStudy (anonymous):

so when it reaches the peak, I just double that number?

OpenStudy (anonymous):

Where will the 6ft come in to play? x0?

OpenStudy (phi):

if we left from height 0. But I would be more careful here and work the equations

OpenStudy (anonymous):

ya, I just got 1.875 and it took it as wrong, so I know I'm not using that 6 right.

OpenStudy (phi):

once we find the max height, we can figure out how long it takes to fall that distance. that will give the time down.

OpenStudy (anonymous):

So I have to find the max height first, and then answer the first question last?

OpenStudy (phi):

yes, at least that is what I'm doing... following my nose on this one.

OpenStudy (phi):

d= x0 + v0 * t - g t^2 x0 is the start distance = 6 v0 = 30 g = 32 t= 15/16

OpenStudy (anonymous):

All right, let me calculate for a second.

OpenStudy (anonymous):

After calculating, I got d=6+28.125-28.125, which would be 6, which was wrong.

OpenStudy (phi):

sorry. what we should do is \[ \int v - g \ t\ dt = v t - \frac{1}{2} g t^2 + c\] in other words we should use d= x0 + v0 * t - (1/2) g t^2

OpenStudy (anonymous):

Recalculating...

OpenStudy (anonymous):

20.0625, it was correct!

OpenStudy (anonymous):

so using that same equation, we find the t?

OpenStudy (phi):

I would just pretend we dropped it from a height of 20.0625 and use d= 1/2 g t^2 and solve for t

OpenStudy (phi):

time up is 0.9375 add time down for total time.

OpenStudy (anonymous):

I got 2.0572795542, which it was saying is wrong...

OpenStudy (phi):

do they ask for you to round ?

OpenStudy (anonymous):

doesn't say so, I tried rounded versions.

OpenStudy (phi):

we can write this equation h(t) = -1/2 g t^2 + 30 t + 6 and solve for when the height is 0, i.e. solve -16 t^2 + 30 t + 6 =0 for t

OpenStudy (anonymous):

Recalculating...

OpenStudy (anonymous):

I got this 1/16 (15+sqrt(321)), for t, and it's a no go on that.

OpenStudy (anonymous):

oh wait a minute, I'm an idiot, hold on.

OpenStudy (anonymous):

oh wait no, I use quadratic and got that answer, what did I do wrong?

OpenStudy (phi):

\[ \frac{1}{2a} ( -b \pm \sqrt{b^2 -4ac}) \]

OpenStudy (anonymous):

ya, used the quadratic formula.

OpenStudy (phi):

but it gives the same t as what you found the good news is h(t) = -16 t^2 + 30 t + 6 has a peak at t= 30/32= 15/16 (using vertex x value is at -b/(2a) ) and h( 15/16) = 20.065 the only problem seems to be the time at which it hits the ground 2.057...

OpenStudy (anonymous):

I'm completely stumped, I'm looking through class notes and don't see anything like this. Great news too, I have another problem just like this after this question for my homework! Looks like I'll be busy for a while.

OpenStudy (anonymous):

Do I need to do something with derivates since that's what section we are in class?

OpenStudy (phi):

we can start with Newton's F= ma and derive the equations we are using but we will get the same answer. what *exactly* have you tried typing in for the total time in the air? they may want just 2 or 2.06 or 2.0573 ?

OpenStudy (anonymous):

2.0573, 2.05, 2.1, 2.06, and then the WHOLE thing.

OpenStudy (phi):

2.1 ?

OpenStudy (anonymous):

ya...rounded up from 2.06?

OpenStudy (anonymous):

It was just a shot in the dark.

OpenStudy (phi):

or we have a typo and they want the time to reach the peak = 15/16

OpenStudy (anonymous):

I have no idea, emailed my professor, hopefully I can get this sorted out. I'll answer back if I get anything. Thanks for the help!

OpenStudy (anonymous):

Emailed him and he says I'm not taking the initial 6 feet into account when I do d= 1/2gt^2

OpenStudy (phi):

if we found the correct max height then we did that part correctly. maybe we catch the rock? so total time is 1.875

OpenStudy (anonymous):

Total time doesn't equal 1.875 according to this, I have 1.875 as total time in the air, and it's giving me a wrong answer.

OpenStudy (phi):

this is also probably wrong, but the only other "total time" would be if we doubled the down time, to get 2.2396 but the total time from leaving the hand 6 ft off the ground at 30 ft/sec until it hits the ground is 2.06 sec anything else is a "bug" or typo. Though if your professor can explain it, I'll retract my assertion.

OpenStudy (anonymous):

Where did you get 2.06?

OpenStudy (phi):

the solution to -16 t^2 + 30 t + 6 =0

OpenStudy (anonymous):

What were you using for t?

OpenStudy (phi):

*** Emailed him and he says I'm not taking the initial 6 feet into account when I do d= 1/2gt^2 **** using that formula would not be correct (though we used it to find the time to drop from the distance of 20.00) we are using -16 t^2 + 30 t + 6 =0 which is the equation of height (or distance) above the ground as a function of time. at time t=0, the height is 6 at time = 15/16 , the height is 20.0625 ft at time = 2.057, the height is 0 ft (i.e. on the ground) if we caught the rock 6 ft above the ground we would solve -16 t^2 + 30 t + 6 = 6 t (-16 t + 30) = 0 --> t=0 or t= 30/16= 1.875 as *possible* value for total time in air. (which you said did not check out)

OpenStudy (phi):

Here is graph of the problem

OpenStudy (anonymous):

.938 is correct! How did you get that?

OpenStudy (anonymous):

So, how do I get the answer to the first problem with this information then?

OpenStudy (phi):

the time from leaving the hand at t=0 until it reaches its max height is t= 15/16 sec 15/16= 0.9375 rounded, 0.938

OpenStudy (phi):

in other words, 0.9375 sec is the time *on the way up* NOT the total time in the air. So we have a typo in the question (always makes it hard to answer when we don't know the question)

OpenStudy (anonymous):

So, all the answers: total time in the air is 2.057, the point that it reaches 20.0625 feet is .938. To get the maximum, we used d= x0 + v0 * t - (1/2) g t^2 x0=6, v0=30, g=32, t=30/32 With that, we got the total time in the air by solving for t of d= 1/2 g t^2, d=20.0625, g = 32. To get at what time it would reach maximum heigh, we divided initial velocity by gravity.

OpenStudy (phi):

looking at the top of the post, it is the first number we found... unfortunately, we thought they wanted the total time in the air, not the time to reach the peak height.

OpenStudy (anonymous):

they wanted both of those, time it took to reach peak height AND the total air time.

OpenStudy (phi):

when you posted .938 is correct! How did you get that? what question was that for?

OpenStudy (anonymous):

For when it would reach the peak height.

OpenStudy (anonymous):

After how many seconds.

OpenStudy (phi):

ok, and did you ever get a correct response for the total time?

OpenStudy (anonymous):

Yeah, it was 2.057, looks like we had the things reversed.

OpenStudy (phi):

we?!

OpenStudy (anonymous):

The next question on the homework now is a general problem about ALL of this. You toss a rock up vertically at an initial speed of V feet per second and release it at an initial height of H feet. The rock will remain in the air for (ANSWER#1) seconds. It will reach a maximum height of (ANSWER#2) feet after (ANSWER#3) seconds.

OpenStudy (anonymous):

hahah, nevermind, I!! I had it reversed, okay?! hahah. Thanks for all your help on that question.

OpenStudy (anonymous):

I already got one answer for the next question. Answer # 3 is V/g

OpenStudy (phi):

for part II, tweak h(t) = -1/2 g t^2 + 30 t + 6 replace 30 with V (initial velocity) and replace 6 with H (initial height) -1/2 g = -1/2 * 32 = -16 so \[ h(t) = -16 t^2 +V t + H \] you can find the peak by find the derivative and setting equal to 0, and then solve for t max height will occur at that t solve for total time by using the quadratic formula on \[ -16 t^2 +V t + H = 0 \] i.e. find the time when height is 0 (rock is on the ground)

OpenStudy (anonymous):

It's being real knitpicky. Won't let me use "t". I'm only allowed to use V, H, and g in my answers =/

OpenStudy (phi):

which question are you answering now?

OpenStudy (anonymous):

Trying to answer all except the last one. I already got that.